7. How much you will have in 15 years if you deposit $1,500 monthly, starting today (at the beginning of the period) in an investment that return 4.75%?
Answer #1

Nper = 15 * 12 = 180
Rate = 4.75%/12 = 0.3958%
Pmt = -1500
PV = 0

Future value can be calculated using the following excel formula:
=FV(rate,nper,pmt,pv,beginning of the period)
=FV(4.75%/12,180,-1500,0,1)
= $394,225.94

Future value = $394,225.94 or $394,226 (Approx).